2-25 Using unit costs for making decisions Rhonda Heninger is a well known software engineer. Her...

2-25

Using unit costs for making decisions

Rhonda Heninger is a well known software engineer. Her specialty is writing software code used in maintaining the security of credit card information. Heninger is approached by the Electronic Commerce Group (ECG). It offers to pay her $120 000 for the right to use her code under licence in its eprocurement software package. Heninger rejects this offer because it provides her with no additional benefits if the eprocurement package is a runaway success. Both parties eventually agree to a contract in which ECG pays Heninger a flat fee of 4120 000 for the right to use her code in up to 10 000 packages. If eprocurement sells more than 10 000 packages, Heninger received 49.60 for each package sold beyond the 10 000 level

  1. What is the unit cost of ECG for Heninger software code included in its eprocurement package if ti sells 2000, 6000, 10000, 20000 packages? Comment results
  2. For prediction fo ECG total cost of using Heninger software code in eprocurement which unit cost of a to a in requirement 1 would you recommend ECG use?
Answer #1

Answer:-

1.         (a) $120,000 ÷ 2,000   = $60.00 per package

(b) $120,000 ÷ 6,000 = $20.00 per package

(c) $120,000 ÷ 10,000 = $12.00 per package

(d) [$120,000 + (10,000 × $9.60)] ÷ 20,000 = $216,000 ÷ 20,000 = $10.80 per package

The unit cost to ECG decreases on a per-unit basis due to the first $120,000 payment being a fixed cost. The $9.60 amount per package beyond 10,000 units is a variable cost. The cost function is:

$360,000 $312,000 $216,000 Total Costs $240,000 $120,000 $120,000 10,000 30,000 20,000 Packages Sold

2.         ECG should not use any of the unit costs in requirement 1 when predicting total costs. Up to 10,000 units, the total cost is a fixed amount.

Beyond 10,000 units, the total cost is a combination of a fixed amount plus a per unit (beyond 10,000 unit) variable amount. The total costs at different volume levels cannot be predicted by using the unit cost at a specific volume level. The total cost should be predicted by combining the total fixed costs and total variable costs rather than by multiplying a unit cost amount by the predicted number of packages sold.